OpenStudy (anonymous):

if the slope is -2 and passes through (2,-9) what is the equation

OpenStudy (anonymous):

Use point-slope form since you have a point of the form (x1,y1) and the slope m. \[y - y1 = m(x - x1)\] where m is the slope, x1 is

OpenStudy (anonymous):

.Use point-slope form since you have a point of the form (x1,y1) and the slope m. y−y1=m(x−x1) where m is the slope, x1 is

OpenStudy (anonymous):

yeah I'm right lol